What is the present worth difference between an investment of $95,000 per year for 45 years and an investment of $95,000 per year forever at an interest rate of 10% per year?

The difference is determined to be $
Answer #1

Answer

PV=A*(P|A,i,n) for finite series

PV=95000*(P|A,10%45)

=95000*9.862807879788514

=936966.749

PV=present value, A=annual investment, i=interest rate=r, n=years

Or

By formula:

PV=$936966.74858

So the value is $936966.74858

=========

for the infinite series

PV=A/i

PV=95000/0.1=950000

============

The difference =950000-936966.749=13033.251